OpenStudy (anonymous):

You have an idea that maybe the time rate of discharge on a capacitor should be proportional to the actual charge on a capacitor. To test this idea, you grab your multi-meter and RC-circuit. So, you power up the capacitor and then switch your circuit off. At that instant (at t=0), you find that the capacitor has 1 coulomb of charge. Ten minutes later, you come back to the circuit and measure 0.89 coulombs (or 89% of the initial charge). Wanting to see how accurate your mathematical model is, you decide that you will predict just how long it will take for the capacitor to reach 1% of its i

OpenStudy (lalaly):

of its what?

OpenStudy (anonymous):

dC/dt = k C(t)

OpenStudy (lalaly):

and what is the question? it is incomplete

OpenStudy (anonymous):

1% of its initial charge. How many minutes does your model predict this will take? whoops

OpenStudy (anonymous):

so solve dC/dt = k C(t), you get dC/C = k dt, integrate both sides ln C = kt + constant C = Ae^(kt)

OpenStudy (anonymous):

when t = 0 , C(t) = .1, so C = e^(kt)

OpenStudy (anonymous):

1*

OpenStudy (lalaly):

\[\frac{dq(t)}{dt}=-k \times q(t)\]it is -k because its discharging \[\frac{dq}{q}=-k dt\]

OpenStudy (anonymous):

ok q(t) = -e^(kt)

OpenStudy (anonymous):

now when t = 10, we have q(10) = .89

OpenStudy (anonymous):

.89 = -e^(k*10) , ...

OpenStudy (anonymous):

wait, that doesnt work.

OpenStudy (anonymous):

i messed up

OpenStudy (anonymous):

q(t) = A e^(-kt)

OpenStudy (anonymous):

you dont have to assume - , since it will come out negative later q(t) = e^(-.011653 t )

OpenStudy (anonymous):

now solve .1 = e^(-.011653 t)

OpenStudy (anonymous):

it will take almost 200 minutes , i get 197.595 minutes

OpenStudy (lalaly):

integrating u get \[\ln(\frac{q}{q_0})=-kt\] q0 is the chaarge when t=0 wer given that when t=10\[\frac{q}{q_0}=0.89\] solve for k then find t when \[\frac{q}{q_0}=0.01\]

OpenStudy (anonymous):

197.59! Thank you! :)

OpenStudy (lalaly):

i got t=395 min:S

OpenStudy (anonymous):

wait,

OpenStudy (anonymous):

i used .1 not .01

OpenStudy (anonymous):

should be 395.18

OpenStudy (lalaly):

yeh thats what i got:D gdjob:D

OpenStudy (anonymous):

lalay i dont get why you wrote q / qo = .89 its ln (q/q0) = .89

OpenStudy (lalaly):

89% of the initial charge so ln(0.89)

OpenStudy (anonymous):

errr, ln ( .89/ q0) = -k (10)

OpenStudy (lalaly):

:)